Solve by completing the square
x^2+4x-3=0

Answer by madhujiithcharu@yahoo.co.in(28)

x^2 + 4x - 3 = 0
x = -4 + or - \/ 16-4(1)(-3) / 2
= -4 + or - \/28 / 2
= -4 + or - 2\/7 /2
= -2 + or - \/7
therefore x = -2 + \/7 , x = -2 - \/7